what point lies on the line with point slope equation y-2=5(x-6)
Answer by ReadingBoosters(3246)

y - 2 = 5(x-6)
Process into slope intercept form
y = 5x - 30 + 2
y = 5x - 28
...
For points, start with the intercepts
..
The y intercept is -28, so (0,-28) is a point on the line
..
The x intercept is where y = 0, so
0 = 5x - 28
5x = 28
x = 28/5
Therefore, (28/5, 0) is another point on the line
..
Choose any value for x and solve for y
x = 15
y = 5(15) - 28 = 75 - 28 = 47
(15,47) is another point on the line
...
and so on
.....................
Have more questions?
Email: HomeworkHelpers@readingboosters.com
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.